OpenStudy (anonymous):

1,3,6,10,15 whats the 15th and 30th term

OpenStudy (campbell_st):

looks like the set of triangular numbers \[a_{n} = a_{n -1} + n\]

OpenStudy (anonymous):

@campbell_st what? im so confused..

OpenStudy (campbell_st):

the sequence is well know.... like the set of square numbers... you can draw a triangle with numbers  the sequence is increasing by the counting numbers1, 1 + 2 = 3, 3 + 3 = 6, 6 + 4 = 10... so the 6th number is 15 + 6 = 21 7th = 21 + 7 = 28 so the formula for a term is \[a_{n} = \frac{n(n + 1)}{2}\] check to see if the formula fits... if it does then its easy to find the 15th and 30th numbers in the sequence